# GSEdit: Efficient Text-Guided Editing of 3D Objects via Gaussian Splatting

### v1

@@ Abstract
Keywords:
Gaussian splatting Radiance fields Inverse rendering 3D Editing
@@
Figure 1: Results of 3D object editing with GSEdit. This method can take as input either meshes or point clouds, and perform an edit in a few minutes, guided by a textual prompt and a diffusion model.
1 Introduction and related work
@@
The fast-paced progress of research in the fields of deep learning and artificial intelligence in recent years brought significant advancement in 3D vision and graphics. Mildenhall et al. [24] were the first to propose the idea of optimizing volumetric color/density functions encoded by deep neural networks (referred to as NeRF, for Neural Radiance Fields) via differentiable volumetric rendering. When the output renders from multiple views are optimized to match ground truth views of some environment, the model changes accordingly to encode the geometry and appearance of the scene.
@@
Due to the exceptional impact it had on the field, NeRF inspired a variety of contributions improving it in several aspects. One of these is rendering time: encoding the color and density functions in a deep and wide MLP, which required millions of queries to render a single view, led to sub-optimal inference performance. The simplest way to approach this problem is to bake the neural functions into efficient discrete structures such as voxel grids [9, 12] or octrees [39, 36]. Alternatively, one could change the rendering model, making the volume integral more efficient [18, 38], or using smaller MLPs in combination with space decomposition [29] or multi-resolution hash encoding [25]. Recently, Kerbl et al. [15] proposed to represent scenes in a hybrid fashion with Gaussian splatting (GS in the following), which employs a finite set of 3D Gaussians to define the continuous color and density fields, while achieving impressive efficiency with an ad-hoc differentiable rasterization algorithm.
@@
Besides a completely new way of approaching the inverse rendering problem, NeRF provided a novel, general framework to handle geometry in an implicit fashion, thus bypassing all issues derived from explicit structures such as triangle meshes. For instance, automated 3D editing/generation tasks were always limited by the complexity of handling changes in the discrete topology, while NeRF-like models allowed to perform these tasks in the neural network’s input/weight space. Two main categories of approaches can be identified: a) using conditional latent codes, either from pretrained image GAN/VAE priors [32, 16, 5, 6, 4, 10] or jointly trained [20, 35], and b) editing/generating the ground truth multi-view images with pretrained diffusion models. This idea was first introduced by Poole et al. in DreamFusion [26], and it rapidly became the state of the art in this line of research due to the popularity and performance of diffusion models. Subsequent work experimented with combining this framework with Instant-NGP [17, 22] as well as training volumetric fields to output latent features to be fed into a Stable Diffusion decoder [23, 30]. DreamGaussian by Tang et al. [34] was the first model for text-to-3D using GS models.
@@
Our work focuses on exploiting the efficiency of GS models in order to define an efficient pipeline for text-guided 3D object editing. Achieving accurate and efficient text-guided editing of 3D objects would have notable implications: mainly, it would mean greater empowerment for 3D artists, allowing them to perform complex, fine grained modifications to their digital environments quickly and automatically. Clearly, this benefit would also similarly reflect on industrial 3D graphics pipelines. Following previous work by Haque et al. [11], we rely on a pretrained instance of the Instruct-Pix2Pix (IP2P) diffusion model [2] to edit ground truth views given a textual prompt, which we use to fine-tune a GS scene encoding the input object. A similar contribution was proposed by Sella et al. [33], however, their voxel-based pipeline requires more than one hour to run to completion. It is worth noting that concurrent work by Fang et al. [8] and Chen et al. [7] also experimented with editing GS scenes via text instructions. Our contributions may be summarized as follows:
@@
  • We adapt the formulation of the Score Distillation Sampling (SDS) loss [26] for GS generation proposed by Tang et al. [34], to GS editing, deriving analytical gradients.
@@
Figure 2: The GSEdit pipeline. 1) It starts by taking a mesh a) or a pretrained Gaussian splatting model b) as input. During this stage, 20 different views of the shape are rendered, and if the input is a mesh, the renders are used to encode it in a GS model. 2) The editing phase consists of picking a camera, rendering the scene from that camera, running a step of Instruct-Pix2Pix [2] to apply the edit, and optimizing the SDS [26] loss. Once the editing is complete, the mesh extraction 3) and texture refinement 4) steps introduced in DreamGaussian [34] are carried out.
2 Background
@@
The idea of representing 3D scenes as Gaussian functions for reconstruction tasks is due to Kerbl et al. [15]. This approach models features (e.g., color) as spatial distribution and employs a fast, differentiable rasterization algorithm to optimize the scene with respect to given ground truth views. Unlike traditional NeRF which requires dense sampling and can be computationally expensive, Gaussian Splatting encodes the scene’s continuous representation via a finite set of Gaussian functions $\{\Theta_{i}\}_{i=1}^{n}$, leading to faster training and rendering times while keeping reconstruction quality similar to NeRF-based methods. In the initial stages, the scene is approximated via an initial guess $\Theta^{0}$ based on 2D projection, which is easy to compute for 3D Gaussians. Each Gaussian within this framework is defined by a covariance matrix, $\Sigma_{i}$ centered around $\mu_{i}$, embodying the spatial distribution as:
$\Theta_{i}(x)=e^{-\frac{1}{2}(x)^{T}\Sigma_{i}^{-1}(x)}$ (1)
@@
$\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}$ (3)
@@
The success of diffusion models in recent years led to a proliferation of text-based image editing/generation methodologies, which span a broad spectrum from fine-tuning practices, as proposed by Kawar et al. [14], to resource-intensive supervised frameworks such as Instruct-Pix2Pix [2].The latter is a diffusion model for image editing based on text conditioning. At its core, the authors created a paired training dataset from the combination of a large language model [3] and a text-to-image model such as Stable Diffusion [30]. This data is used to train a conditional Diffusion model that outputs an edited image from a source image and a user text prompt. IP2P operates on the latent space since studies have demonstrated that efficiency and quality are improved in Latent Diffusion Models [1]. Given an input image $x$, a noise $\epsilon\sim\mathcal{N}(0,1)$, scaled according to the timestamp $t$, is added to the latent representation of the input $z=E(x)$ to produce a noisy latent $z_{t}$. From here a U-Net $\epsilon_{\phi}$ is trained to predict the noise of $z_{t}$ given the input prompt $c_{T}$ and the conditioned image $c_{I}$ with the aim of minimizing the loss:
$\left.L=\mathbb{E}_{E(x),E\left(c_{I}\right),c_{T},\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\phi}\left(z_{t},t,E\left(c_{I}\right),c_{T}\right)\right)\|_{2}^{2}\right]$ (4)
The model employs a U-Net architecture [31], renowned for its efficacy in image-to-image translation tasks, to modulate edits as instructed by text embeddings. These embeddings, derived from user-provided textual instructions, guide the model to perform contextually relevant edits while maintaining coherence with the source image. The interaction between the image condition and text embeddings within the U-Net framework enables precise, instruction-aligned modifications, highlighting the model’s ability to intuitively adapt to diverse editing tasks.
Refer to captionInput Refer to captionEdit Refer to captionInput Refer to captionEdit
‘‘Turn it into a wolf’’ ‘‘Make it look realistic’’
Refer to captionInput Refer to captionEdit Refer to captionInput Refer to captionEdit
‘‘Turn it into a robot’’ ‘‘Make it made of gold’’
Figure 3: A collection of input/output pairs for our GSEdit pipeline. The results show that our method is flexible enough to handle various settings similarly well, such as changes of object identity (dog $\rightarrow$ wolf, gnome $\rightarrow$ robot) and visual style (realistic duck, golden ice-cream). GSEdit aims to perform the editing while preserving the input’s features, such as pose, overall shape, or the hat and lantern of the gnome.
3 Method
@@
3.1 Input Representation
In order to keep our method as general as possible, we implemented it to support any Gaussian splatting scene as input. However, for the scope of this paper, we are interested in working with single foreground objects; therefore, we consider the following cases as valid inputs for our pipeline: a) GS reconstructions, computed on-the-fly from multi-view renders of a given textured 3D mesh. Foreground/background segmentation in this “synthetic” case is trivial, and Gaussian splatting optimization always manages to reconstruct the foreground object alone, without background “ghostly artifacts” [37] which were common for previous NeRF-like models. b) GS scenes output by generative models such as DreamGaussian [34]. We chose DreamGaussian as a generation procedure for part of our experiment scenes due to this model being tailored for 3D object generation, rather than complete scenes. Given a textual prompt, DreamGaussian rapidly generates 3D Gaussians matching the instructions. We provide examples of such inputs for our pipeline in Figure 2.
3.2 3D Object Editing
For the editing part, we use the same cameras placed for the generation of the Gaussian Splatting to retrieve in each training step the current render of the scene [^1] If the input is generated by DreamGaussian, we place random cameras and generate multi-view renders accordingly, as we do in the case of training a GS from scratch.. At each step, we sample a random camera $p$ orbiting the object center and render the RGB image $I^{p}_{RGB}$. This render is subsequently processed through the IP2P model along with the input textual prompt for editing. We adopt a singular step adjustment strategy akin to the one employed in Dream Gaussian [34], focusing on incremental and directional enhancements within the CLIP space to preserve the integrity of the scene. Once all of the rendered images are processed, the parameters of the Gaussians are updated via differentiable rendering to match the current ground truth renders. Afterward, images to feed into the IP2P model are updated by rendering the new GS scene. This cyclical process of capturing, editing, and updating methodically evolves the original scene, allowing a gradual and controlled transition until the editing state has converged (i.e., IP2P does not apply significant modifications to the current state).
@@ 3.1 Input Representation
We now explain how the Score Distillation Loss [26] (SDS) can be adapted to GS editing. Figure 4 summarizes this process: given a timestep $t$, a noise $\epsilon\sim\mathcal{N}(0,I)$ (2) is added to the $i$-th rendered image encoded with IP2P (1), resulting in a noisy image $z_{t}$ (3). Afterward, the pretrained IP2P U-Net $\epsilon_{\phi}$ tries to predict the amount of noise $\hat{\epsilon}$ (4) in the input image $z_{t}$ given the original image $c_{I}$ and the text embeddings for the editing instruction $c_{T}$:
$\hat{\epsilon}=\epsilon_{\phi}(z_{t};t,c_{I},c_{T})$ (5)
Figure 4: A visualization of how the edits performed by the U-Net of IP2P are exploited to guide the edit.
@@
$\Delta_{i}=\hat{\epsilon}-\epsilon$ (6)
@@
$\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t\sim\mathcal{U}(1,T)}\left[w(t)\Delta_{i}\dfrac{\partial z_{t}}{\partial\Theta}\right]$ (7)
where $w(t)$ is a weighting function that depends on the timestep $t$. We clip $t$ in a $[t_{min},t_{max}]$ range and we linearly decrease this value as in Huang et al. [13] so that the amount of noise added to the latent render $z_{i}$ is greater in the initial steps and decreases over time.
3.3 Mesh Extraction and Texture Refinement
To efficiently extract and refine a polygonal mesh from the edited point cloud, we leverage the method introduced in [34]. Their method is divided into a first part that extracts the mesh, and a second part that refines textures. The former consists of two steps:
  1. 1.
    Local Density Query. To extract the mesh geometry, they rely on Marching Cubes [21]. The first step is to reduce the number of Gaussians by dividing the space into $16^{3}$ and removing all those Gaussians whose center is located outside each local block. Then, they query a $8^{3}$ dense grid inside each block and sum up the weighted opacity of each of the $k$ Gaussian in the block:
    $d(x)=\sum_{i=1}^{k}\alpha_{i}\exp\left(-\frac{1}{2}(x-x_{i})^{T}\Sigma_{i}^{-1}(x-x_{i})\right)$ (8)
@@ 3.3 Mesh Extraction and Texture Refinement
$I^{p}_{\text{fine}}=f_{\phi}(I^{p}_{\text{coarse}}+\epsilon(t_{\text{start}});t_{\text{start}},\hat{c}_{T})$ (9)
@@
$\mathcal{L}_{\text{MSE}}=\lVert I^{p}_{\text{fine}}-I^{p}_{\text{coarse}}\rVert^{2}$ (10)
Figure 5: Examples of applications of GSEdit with fixed input mesh, varying the editing instruction. The results show the ability of our model to change object identity while keeping original features (cactus), adding simple features (hat), editing “artistic” style (lowpoly), and material (marble).
3.4 Implementation Details

### v2

@@ Abstract.
Gaussian splatting and Radiance fields and Inverse rendering and 3D Editing
Gaussian splatting, Radiance fields, Inverse rendering, 3D Editing
ccs: Computing methodologies Shape modelingccs: Computing methodologies Shape representationsccs: Computing methodologies Appearance and texture representationsccs: Computing methodologies Rendering
@@
Figure 1. Results of 3D object editing with GSEdit. This method can take as input either meshes or point clouds, and perform an edit in a few minutes, guided by a textual prompt and a diffusion model.
1. Introduction and related work
Figure 2. The GSEdit pipeline. 1) It starts by taking a mesh a) or a pretrained Gaussian splatting model b) as input. During this stage, 20 different views of the shape are rendered, and if the input is a mesh, the renders are used to encode it in a GS model. 2) The editing phase consists of picking a camera, rendering the scene from that camera, running a step of Instruct-Pix2Pix (Brooks et al., 2023) to apply the edit, and optimizing the SDS (Poole et al., 2023) loss. Once the editing is complete, the mesh extraction 3) and texture refinement 4) steps introduced in DreamGaussian (Tang et al., 2023) are carried out.
@@ 1. Introduction and related work
The fast-paced progress of research in the fields of deep learning and artificial intelligence in recent years brought significant advancement in 3D vision and graphics. Mildenhall et al. (2021) were the first to propose the idea of optimizing volumetric color/density functions encoded by deep neural networks (referred to as NeRF, for Neural Radiance Fields) via differentiable volumetric rendering. When the output renders from multiple views are optimized to match ground truth views of some environment, the model changes accordingly to encode the geometry and appearance of the scene.
@@
Due to the exceptional impact it had on the field, NeRF inspired a variety of contributions improving it in several aspects. One of these is rendering time: encoding the color and density functions in a deep and wide MLP, which required millions of queries to render a single view, led to sub-optimal inference performance. The simplest way to approach this problem is to bake the neural functions into efficient discrete structures such as voxel grids (Garbin et al., 2021; Hedman et al., 2021) or octrees (Yu et al., 2021; Wang et al., 2022). Alternatively, one could change the rendering model, making the volume integral more efficient (Lindell et al., 2021; Wu et al., 2021), or using smaller MLPs in combination with space decomposition (Reiser et al., 2021) or multi-resolution hash encoding (Müller et al., 2022). Recently, Kerbl et al. (2023) proposed to represent scenes in a hybrid fashion with Gaussian splatting (GS in the following), which employs a finite set of 3D Gaussians to define the continuous color and density fields, while achieving impressive efficiency with an ad-hoc differentiable rasterization algorithm.
@@
Besides a completely new way of approaching the inverse rendering problem, NeRF provided a novel, general framework to handle geometry in an implicit fashion, thus bypassing all issues derived from explicit structures such as triangle meshes. For instance, automated 3D editing/generation tasks were always limited by the complexity of handling changes in the discrete topology, while NeRF-like models allowed to perform these tasks in the neural network’s input/weight space. Two main categories of approaches can be identified: a) using conditional latent codes, either from pretrained image GAN/VAE priors (Schwarz et al., 2020; Kosiorek et al., 2021; Chan et al., 2021, 2022; Cai et al., 2022; Gu et al., 2022) or jointly trained (Liu et al., 2021; Wang et al., 2021), and b) editing/generating the ground truth multi-view images with pretrained diffusion models. This idea was first introduced in DreamFusion by Poole et al. (2023), and it rapidly became the state of the art in this line of research due to the popularity and performance of diffusion models. Subsequent work experimented with combining this framework with Instant-NGP (Lin et al., 2023; Melas-Kyriazi et al., 2023) as well as training volumetric fields to output latent features to be fed into a Stable Diffusion decoder (Metzer et al., 2023; Rombach et al., 2022). DreamGaussian (Tang et al., 2023) was the first model for text-to-3D using GS models.
@@
Our work focuses on exploiting the efficiency of GS models in order to define an efficient pipeline for text-guided 3D object editing. Achieving accurate and efficient text-guided editing of 3D objects would have notable implications: mainly, it would mean greater empowerment for 3D artists, allowing them to perform complex, fine grained modifications to their digital environments quickly and automatically. Clearly, this benefit would also similarly reflect on industrial 3D graphics pipelines. Following previous work by Haque et al. (2023), we rely on a pretrained instance of the Instruct-Pix2Pix (IP2P) diffusion model (Brooks et al., 2023) to edit ground truth views given a textual prompt, which we use to fine-tune a GS scene encoding the input object. A similar contribution was proposed by Sella et al. (2023), however, their voxel-based pipeline requires around one hour to run to completion. GS models were previously applied to 3D scene editing by Fang et al. (2024) and Chen et al. (2023), with a different focus than our work. Our contributions may be summarized as follows:
@@
  • We adapt the formulation of the Score Distillation Sampling (SDS) loss (Poole et al., 2023) for GS generation proposed by Tang et al. (2023), to GS editing, deriving analytical gradients.
@@
2. Background
@@
The idea of representing 3D scenes as Gaussian functions for reconstruction tasks is due to Kerbl et al. (2023). This approach models features (e.g., color) as spatial distribution and employs a fast, differentiable rasterization algorithm to optimize the scene with respect to given ground truth views. Unlike traditional NeRF which requires dense sampling and can be computationally expensive, Gaussian Splatting encodes the scene’s continuous representation via a finite set of Gaussian functions $\{\Theta_{i}\}_{i=1}^{n}$, leading to faster training and rendering times while keeping reconstruction quality similar to NeRF-based methods. In the initial stages, the scene is approximated via an initial guess $\Theta^{0}$ based on 2D projection, which is easy to compute for 3D Gaussians. Each Gaussian within this framework is defined by a covariance matrix, $\Sigma_{i}$ centered around $\mu_{i}$, embodying the spatial distribution as:
(1) $\Theta_{i}(x)=e^{-\frac{1}{2}(x)^{T}\Sigma_{i}^{-1}(x)}$
@@
(3) $\mathcal{L}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{D-SSIM}$
@@
The success of diffusion models in recent years led to a proliferation of text-based image editing/generation methodologies, which span a broad spectrum from fine-tuning practices, as proposed by Kawar et al. (2023), to resource-intensive supervised frameworks such as Instruct-Pix2Pix (Brooks et al., 2023).The latter is a diffusion model for image editing based on text conditioning. At its core, the authors created a paired training dataset from the combination of a large language model (Brown et al., 2020) and a text-to-image model such as Stable Diffusion (Rombach et al., 2022). This data is used to train a conditional Diffusion model that outputs an edited image from a source image and a user text prompt. IP2P operates on the latent space since studies have demonstrated that efficiency and quality are improved in Latent Diffusion Models (Blattmann et al., 2022). Given an input image $x$, a noise $\epsilon\sim\mathcal{N}(0,1)$, scaled according to the timestamp $t$, is added to the latent representation of the input $z=E(x)$ to produce a noisy latent $z_{t}$. From here a U-Net $\epsilon_{\phi}$ is trained to predict the noise of $z_{t}$ given the input prompt $c_{T}$ and the conditioned image $c_{I}$ with the aim of minimizing the loss:
(4) $L=\mathbb{E}_{E(x),E\left(c_{I}\right),c_{T},\epsilon\sim\mathcal{N}(0,1),t}\left[\left\lVert\epsilon-\epsilon_{\phi}\left(z_{t},t,E\left(c_{I}\right),c_{T}\right)\right\rVert_{2}^{2}\right]$
The model employs a U-Net architecture (Ronneberger et al., 2015), renowned for its efficacy in image-to-image translation tasks, to modulate edits as instructed by text embeddings. These embeddings, derived from user-provided textual instructions, guide the model to perform contextually relevant edits while maintaining coherence with the source image. The interaction between the image condition and text embeddings within the U-Net framework enables precise, instruction-aligned modifications, highlighting the model’s ability to intuitively adapt to diverse editing tasks.
Refer to captionInput Refer to captionEdit
‘‘Turn it into a wolf’’
Refer to captionInput Refer to captionEdit
‘‘Make it look realistic’’
Refer to captionInput Refer to captionEdit
‘‘Turn it into a robot’’
Refer to captionInput Refer to captionEdit
‘‘Make it made of gold’’
Figure 3. A collection of input/output pairs for our GSEdit pipeline. The results show that our method is flexible enough to handle various settings similarly well, such as changes of object identity (dog $\rightarrow$ wolf, gnome $\rightarrow$ robot) and visual style (realistic duck, golden ice-cream). GSEdit aims to perform the editing while preserving the input’s features, such as pose, overall shape, or the hat and lantern of the gnome.
3. Method
@@
3.1. Input Representation
In order to keep our method as general as possible, we implemented it to support any Gaussian splatting scene as input. However, for the scope of this paper, we are interested in working with single foreground objects; therefore, we consider the following cases as valid inputs for our pipeline: a) GS reconstructions, computed on-the-fly from multi-view renders of a given textured 3D mesh. Foreground/background segmentation in this “synthetic” case is trivial, and Gaussian splatting optimization always manages to reconstruct the foreground object alone, without background “ghostly artifacts” (Warburg* et al., 2023) which were common for previous NeRF-like models. b) GS scenes output by generative models such as DreamGaussian (Tang et al., 2023). We chose DreamGaussian as a generation procedure for part of our experiment scenes due to this model being tailored for 3D object generation, rather than complete scenes. Given a textual prompt, DreamGaussian rapidly generates 3D Gaussians matching the instructions. We provide examples of such inputs for our pipeline in Figure 2.
3.2. 3D Object Editing
For the editing part, we use the same cameras placed for the generation of the Gaussian Splatting to retrieve in each training step the current render of the scene [^1] If the input is generated by DreamGaussian, we place random cameras and generate multi-view renders accordingly, as we do in the case of training a GS from scratch.. At each step, we sample a random camera $p$ orbiting the object center and render the RGB image $I^{p}_{RGB}$. This render is subsequently processed through the IP2P model along with the input textual prompt for editing. We adopt a singular step adjustment strategy akin to the one employed in Dream Gaussian (Tang et al., 2023), focusing on incremental and directional enhancements within the CLIP space to preserve the integrity of the scene. Once all of the rendered images are processed, the parameters of the Gaussians are updated via differentiable rendering to match the current ground truth renders. Afterward, images to feed into the IP2P model are updated by rendering the new GS scene. This cyclical process of capturing, editing, and updating methodically evolves the original scene, allowing a gradual and controlled transition until the editing state has converged (i.e., IP2P does not apply significant modifications to the current state).
Figure 4. Examples of applications of GSEdit with fixed input mesh, varying the editing instruction. The results show the ability of our model to change object identity while keeping original features (cactus), adding simple features (hat), editing “artistic” style (lowpoly), and material (marble).
@@ 3.1. Input Representation
We now explain how the Score Distillation Loss (Poole et al., 2023) (SDS) can be adapted to GS editing. Figure 5 summarizes this process: given a timestep $t$, a noise $\epsilon\sim\mathcal{N}(0,I)$ (2) is added to the $i$-th rendered image encoded with IP2P (1), resulting in a noisy image $z_{t}$ (3). Afterward, the pretrained IP2P U-Net $\epsilon_{\phi}$ tries to predict the amount of noise $\hat{\epsilon}$ (4) in the input image $z_{t}$ given the original image $c_{I}$ and the text embeddings for the editing instruction $c_{T}$:
(5) $\hat{\epsilon}=\epsilon_{\phi}(z_{t};t,c_{I},c_{T})$
Figure 5. A visualization of how the edits performed by the U-Net of IP2P are exploited to guide the edit.
@@
(6) $\Delta_{i}=\hat{\epsilon}-\epsilon$
@@
(7) $\nabla_{\Theta}\mathcal{L}_{\text{SDS}}=\mathbb{E}_{t\sim\mathcal{U}(1,T)}\left[w(t)\Delta_{i}\dfrac{\partial z_{t}}{\partial\Theta}\right]$
where $w(t)$ is a weighting function that depends on the timestep $t$. We clip $t$ in a $[t_{min},t_{max}]$ range and we linearly decrease this value as proposed by Huang et al. (2023) so that the amount of noise added to the latent render $z_{i}$ is greater in the initial steps and decreases over time.
3.3. Mesh Extraction and Texture Refinement
To efficiently extract and refine a polygonal mesh from the edited point cloud, we leverage the method introduced in (Tang et al., 2023). Their method is divided into a first part that extracts the mesh, and a second part that refines textures. The former consists of two steps:
  1. (1)
    Local Density Query. To extract the mesh geometry, they rely on Marching Cubes (Lorensen and Cline, 1987). The first step is to reduce the number of Gaussians by dividing the space into $16^{3}$ and removing all those Gaussians whose center is located outside each local block. Then, they query a $8^{3}$ dense grid inside each block and sum up the weighted opacity of each of the $k$ Gaussian in the block:
    (8) $d(x)=\sum_{i=1}^{k}\alpha_{i}\exp\left(-\frac{1}{2}(x-x_{i})^{T}\Sigma_{i}^{-1}(x-x_{i})\right)$
@@ 3.3. Mesh Extraction and Texture Refinement
(9) $I^{p}_{\text{fine}}=f_{\phi}\left(I^{p}_{\text{coarse}}+\epsilon(t_{\text{start}});t_{\text{start}},\hat{c}_{T}\right)$
@@
(10) $\mathcal{L}_{\text{MSE}}=\left\lVert I^{p}_{\text{fine}}-I^{p}_{\text{coarse}}\right\rVert^{2}$
Figure 6. Multi-view renders of three meshes edited with GSEdit. The results show good consistency of the edited shapes with respect to the view direction.
Input shape Prompt IN2N Vox-E GaussianEditor GSEdit (ours)
Refer to caption Refer to caption Refer to $\sim$23 min Refer to $\sim$50 min Refer to $\sim$6.5 min Refer to $\sim$4 min
Refer to caption Refer to caption Refer to $\sim$20 min Refer to $\sim$50 min Refer to $\sim$6.5 min Refer to $\sim$4 min
Refer to caption Refer to caption Refer to $\sim$13 min Refer to $\sim$50 min Refer to $\sim$6.5 min Refer to $\sim$4 min
Refer to caption Refer to caption Refer to $\sim$41 min Refer to $\sim$50 min Refer to $\sim$6.5 min Refer to $\sim$4 min
Figure 7. Qualitative comparison of results for object editing. The outputs of Instruct-NeRF2NeRF Haque et al. (2023) show its inability to focus detail on a single object, hinting at the method only being suitable for entire scenes. Despite the additional contributions, the same weakness is partially inherited by GaussianEditor (Chen et al., 2023) due to employing the same backbone architecture. The results of Vox-E (Sella et al., 2023) can be more detailed than ours in some settings, however at a much higher computational cost and with no obvious way of extracting the geometry as a 3D mesh.
3.4. Implementation Details